OpenStudy (jenniferjuice):

What is the equation of the quadratic graph with a focus of (3, 4) and a directrix of y = 8? f(x) = −one eighth (x − 3)2 + 4 f(x) = −one eighth (x − 3)2 + 6 f(x) = one eighth (x − 3)2 + 4 f(x) = one eighth (x − 3)2 + 6

OpenStudy (jenniferjuice):

@IMStuck non free medal here ^.^

OpenStudy (imstuck):

lol!!!!!!!

OpenStudy (imstuck):

Center is at (3,6)

OpenStudy (imstuck):

And this is a negative parabola, meaning it opens downward, becuase the focus is below the directrix. SO...

OpenStudy (jenniferjuice):

Sooooooooo

OpenStudy (imstuck):

It's the second choice down.

OpenStudy (imstuck):

y = 4px^2 and p = -1/2

OpenStudy (imstuck):

4p=-2 so p = -1/2. Do you see that part? that's the tricky part of this

OpenStudy (jenniferjuice):

i see it

OpenStudy (imstuck):

TY for the non-free medal! Been a pleasure! I love these things, as ridiculously tedious and difficult as they are!

OpenStudy (jenniferjuice):

please help with the other question i tagged you in!?

OpenStudy (imstuck):

p is the number of units that the focus is away from the center. The focus is down 2 units so x = -2...4p=-2 and p = -1/2